### v3

# Social Bias in Large Language Models For Bangla: An Empirical Study on Gender and Religious Bias

Abstract
The rapid growth of Large Language Models (LLMs) has put forward the study of biases as a crucial field. It is important to assess the influence of different types of biases embedded in LLMs to ensure fair use in sensitive fields. Although there have been extensive works on bias assessment in English, such efforts are rare and scarce for a major language like Bangla. In this work, we examine two types of social biases in LLM generated outputs for Bangla language. Our main contributions in this work are: (1) bias studies on two different social biases for Bangla, (2) a curated dataset for bias measurement benchmarking and (3) testing two different probing techniques for bias detection in the context of Bangla. This is the first work of such kind involving bias assessment of LLMs for Bangla to the best of our knowledge. All our code and resources are publicly available for the progress of bias related research in Bangla NLP.  [^1] https://github.com/csebuetnlp/BanglaSocialBias
Social Bias in Large Language Models For Bangla: An Empirical Study on Gender and Religious Bias
Jayanta Sadhu, Maneesha Rani Saha, Rifat Shahriyar Bangladesh University of Engineering and Technology (BUET) {1705047, 1805076}@ugrad.cse.buet.ac.bd, rifat@cse.buet.ac.bd
1 Introduction
The rapid advancement of Large Language Models (LLMs) has significantly impacted various domains, particularly in social influence and the technology industry Kasneci et al. (2023); Dong et al. (2024b). Given their growing influence, it is crucial to ensure LLMs are free from harmful biases to avoid legal and ethical issues Weidinger et al. (2022); Deshpande et al. (2023). In the context of computing/socio-technical systems, bias refers to the unfair and systematic favoritism shown towards certain individuals or social groups, often at the expense of others, resulting in discriminatory outcomes Friedman and Nissenbaum (1996); Blodgett et al. (2020). Hence, analyzing bias and stereotypical behavior in LLMs is vital for identifying and mitigating existing biases.
Bangla, the sixth most spoken language globally with over 230 million native speakers constituting 3% of the world’s population [^2] https://w.wiki/Psq, has remained underrepresented in NLP literature due to a lack of quality datasets (Joshi et al., 2020). This gap limits our understanding of bias characteristics in language models, including LLMs. Historically, societal views in Bangla-speaking regions have undervalued women, leading to employment and opportunity discrimination (Jain et al., 2021; Tarannum, 2019). Additionally, the region’s cultural and historical context between two major religions, Hindu and Muslim, makes Bangla a valuable case study for examining religious biases as well.
In this study, we pose the question, to what extent do multilingual LLMs exhibit Gender and Religious Bias in Bangla context?. To address this, we present: (1) a curated dataset specifically designed to detect gender and religious biases in Bangla, (2) detailed bias probing analysis on both popular and state-of-the-art closed and open-source LLMs, and (3) an empirical study on bias through LLM-generated responses.
Our findings reveal significant biases in LLMs for the Bangla language and highlight shortcomings in their generative power and understanding of the language, underscoring the need for future de-biasing efforts and better Bangla specific finetuning of LLMs.
Figure 1: Workflow for the creation of naturally sourced corpus for the experiments detailed in this study.
2 Related Work
Existence of gender bias has been exposed in tasks like Natural Language Understanding Bolukbasi et al. (2016); Gupta et al. (2022); Stanczak and Augenstein (2021) and Natural Language Generation Sheng et al. (2019); Lucy and Bamman (2021); Huang et al. (2021). Benchmarks such as WinoBias Zhao et al. (2018) and Winogender Rudinger et al. (2018) have been used to measure gender biases in LMs. Preliminary studies on religious and ethnic biases are done in some works (BehnamGhader and Milios, 2022; Navigli et al., 2023; Abid et al., 2021). Works like Nadeem et al. (2021); Nangia et al. (2020) provide frameworks and datasets for different types of biases including gender and religion. IndiBias Sahoo et al. (2024), a benchmark in Indian context, has been introduced to measure socio-cultural biases in LLMs.
Recent studies have conducted experiments on determining gender stereotypes in LLMs Kotek et al. (2023); Ranaldi et al. (2024); Jha et al. (2023); Dong et al. (2024a) and debiasing techniques Gallegos et al. (2024); Ranaldi et al. (2024), but most of them are on English. There are a few works on multilingual settings Zhao et al. (2024a); Vashishtha et al. (2023), but such efforts are not common for Bangla. The most preliminary work on Bangla bias detection is found in the works of Sadhu et al. (2024), that includes static and contextual embeddings. Effectiveness of varied probing techniques for extracting cultural variations in pretrained LMs has been discussed in Arora et al. (2023).
3 Linguistic Characteristics of Bangla Pronouns
Unlike English and similar languages, Bangla lacks gender-specific pronouns (e.g., he, she). Instead, Bangla employs common pronouns that are used interchangeably for both male and female genders in both singular and plural forms. Moreover, the structure of Bangla sentences does not change in terms of verbs or other grammatical elements to indicate the gender of the subject, as is the case in languages like Hindi or Spanish. As a result, sentences in Bangla that do not include gender-specific nouns or proper names are inherently gender-neutral.
4 Data
We use two strategies for LLM probing: Template Based and Naturally Sourced. The template-based approach uses curated templates for gendered persona or religious group predictions for bias evaluation. Naturally sourced sentences, on the other hand, are used to make explicit predictions about groups or genders, helping to understand the LLM’s ability to interpret natural scenarios. We explain the two techniques as follows:
Template Based: We create semantically bleached templates with placeholders for specific traits, filled with adjective words from categories like Personality, Outlook, Communal, and Occupation (see Figures 6 and 9 in appendix). The adjective categories and words were validated by native Bangla-speaking authors. To explore the effect of occupation on role prediction, we intermix professions with traits in the templates. Examples in the Placeholder column of Figure 9 illustrate the process. Care was taken to avoid stereotypes, ensuring all adjectives and occupations were equally probable for any gender or religious community. For gender detection, the templates employed gender-neutral pronouns of Bangla, along with simple and context-independent sentences to obscure any clues about the gender of the person being referred to. Similarly, for detecting bias related to religious communities, the templates used common, non-specific pronouns (e.g., they/them) and avoided any contextual or identifying details that could hint at the religious affiliation of the individual mentioned in the prompt. In total, we have 2772 template sentences by combining both the categories (see Appendix 4 for detailed statistics).
Figure 2: Workflow of Filtering Naturally Sourced Data using LLM and Prompt Preparation
Naturally Sourced: The workflow of preparing the corpus for naturally sourced sentences is illustrated in Figure 1. We use the BIBED dataset Das et al. (2023), specifically the Explicit Bias Evaluation (EBE) data for naturally occurring scenarios. The sentences are structured in pairs, each containing one identifying subject from a group of either male-female words (for gender) or Hindu-Muslim words (for religion). Figure 7 (in the appendix) illustrates how sentences are grouped into ’Gender’ and ’Religion’ biases. It provides original (root) sentences, paired sentences with altered gender or religion entities, and the modifications necessary to transform them into data points.
An important limitation of the BIBED dataset is that many sentences are not equally probable for both contrasting identities due to issues such as contradictory historical facts, entity-specific information not applicable to the other, incorrect identification of gender or religion entity in the root sentences, or lack of moderation. Examples of these non-applicable scenarios are shown in Figure 8 (in Appendix). To address this, we manually curated sentences to ensure equal applicability to both identities (see Appendix C for details). Each selected root sentence was transformed into a data point by removing the main identifying subject (male-female for gender or Hindu-Muslim for religion) and converting it into a bias detection prompt. Examples of the final prompt format are provided in the Modification column of Figure 7. The prompt creation workflow is illustrated in Figure 2. After curation, 2416 pairs were retained for gender and 1535 for religion.
5 Experimental Setup
5.1 Model Selection
For our experiment we provide results for four state-of-the-art LLMs: Llama3-8b (version: Meta-Llama-3-8B-Instruct 333meta-llama/Meta-Llama-3-8B-Instruct) (AI@Meta, 2024), GPT-3.5-Turbo 444gpt-3-5-turbo, GPT-4o 555gpt-4o and Claude-3.5-Sonnet666anthropic/claude-3.5-sonnet. To reduce randomness, we set the temperature very low ($temp=0.1$) and restrict the maximum response length to 128. Since Bangla is a low resource language, not many models could generate the expected response we required. Some of the open source models that we used but failed to get presentable results are mentioned in the limitations section.
5.2 Prompt
In the case of template based probing, we prompt the model for gendered role or religious identity selection, and in the case of naturally sourced probing, we use fill in the blanks approach.
Template Probing: As shown in Table 5 (appendix F), LLMs are instructed to respond with a gender or religion assuming role of a Bengali person for template based probing. Each input contains a sentence with gender neutral pronoun along with one of the trait words listed in Figure 6. Input sentence templates with placeholders are explained in Figure 9.
Naturally Sourced Probing: LLMs are instructed to fill in the blank with a gender (male-female) or religion (Hindu-Muslim) reflecting the context of the input. Modification of EBE datapoints for prompt creation is shown in Figure 7.
In table 1, we provide the number of unique prompts for each model.
Probing Method Category # Prompts
Template Based Gender 2128
Religion 644
Naturally Sourced Gender 2416
Religion 1535
Table 1: Probing Methods, Categories, and Number of Prompts for each LLM
During evaluation, the options (gender or religion prediction) provided to LLMs inside a prompt are randomly shuffled for both gender and religious entities to avoid selection bias Zheng et al. (2024).
5.3 Evaluation Metric
We employ the widely used fairness metric, Disparate Impact (DI) Feldman et al. (2015), calculated as $\frac{P(Y=1|S\neq 1)}{P(Y=1|S=1)}$. For our binary identifiers (e.g., male-female, Hindu-Muslim), DI can be applied through empirical estimation. In task Q, for category a with outcomes x and y, DI is calculated by the following formula:
$DI_{Q}(a)=\frac{P(Q=x|a)}{P(Q=y|a)}$
We use occurrence frequency instead of probability Zhao et al. (2024b) and adjust the metric to adjust equal proportionality in bias scores (further justification and detail is provided in appendix B):
$Bias\>Score=DI_{Q}(a)=\tanh\left(\log\frac{C_{x}(a)}{C_{y}(a)}\right)$
Here, $C_{z}$ represents the frequency of class z. We compute $DI_{G}$ and $DI_{R}$ for gender and religion biases, where $(x=female,y=male)$ and $(x=Hindu,y=Muslim)$. For a fair LLM, the DI score should be close to 0.
5.4 Metric Interpretation and Bias Direction
To better understand the bias score from numerical values, we provide an interpretation framework in Table 2. Greater deviation from the neutral line denotes the presence of greater bias in either directions.
Bias Type Bias Score
Positive Negative
Gender Female-biased Male-biased
Religion Hindu-biased Muslim-biased
Table 2: Interpretation of Bias Scores for Gender and Religion
6 Results and Evaluation
(a) Bias Scores for Gender Bias (Positive Traits)
(b) Bias Scores for Gender Bias (Negative Traits)
(c) Bias Scores for Religious Bias (Positive Traits)
(d) Bias Scores for Religious Bias (Negative Traits)
Figure 3: Bias Scores in role selection for multiple LLMs in the case of template based probing for gender and religion data. Positive and negative traits results are shown separately. The neutral line $(Bias\>Score=0)$ is highlighted in all the figures. The positive bias scores in figures 3(a) and 3(b) represents Female biased and in figures 3(c) and 3(d) represents Hindu biased. (Note that the results for Occupation are the same for positive and negative traits and only included in contrasting graphs for the ease of comprehending the effect of inter-mixing with other traits.)
6.1 Template Based Probing Results
We present the template based results in figure 3. We report the results based on seven different categories and include the results for positive and negative traits separately for more nuanced variations.
Gender Bias: Our findings (Figure 3(a), 3(b)) show that GPT-3.5-Turbo is consistently biased toward females, while Llama-3 and Claude-3.5-Sonnet are biased toward males across both positive and negative traits. GPT-4o exhibits the most fluctuation, switching its bias depending on the category. When the traits change from positive to negative, GPT-4o changes substantially from female direction to male direction for Personality and Communal based traits. Except for GPT-3.5-Turbo, all models display a strong male bias for occupations.
Inclusion of occupation in prompts had the most significant impact on GPT-4o, reversing its bias direction. In most other cases, occupations shifted bias scores further towards males, suggesting that LLMs place significant weight on occupation when inferring gender. High negative bias scores of Claude-3.5-Sonnet, compared to other models, may be due to the limitations in understanding Bangla context, warranting further investigation.
Religious Bias: For positive traits (Figure 3(c)), all the LLMs exhibit positive bias scores, i.e. being biased for Hindu Religion followers. All LLMs show positive scores for Occupation. The responses form GPT-4o and Llama-3 hold neutral positions for Outlook, but when associated with Occupation, their position of neutrality is compromised. For Llama-3, no specific pattern is evident and high fluctuations are noticeable.
For negative traits (Figure 3(d)), GPT models tend to adopt a neutral stance when Outlook adjectives are included in prompts. We hypothesize that the models avoid offensive responses by maintaining neutrality in negative contexts. However, GPT-4o shows a significant bias towards Muslims when negative ideological elements are present, which is concerning.
6.2 Naturally Sourced Probing Results
Figure 4: Bias results in Naturally Sourced(EBE) probing method for multiple LLMs
Gender Bias: Figure 4 shows that GPT-4o has the highest bias score, indicating a significant gender disparity in its performance. GPT-3.5, with a score just above neutral, demonstrates relatively balanced results with minor disparities. Llama-3, with a negative bias score, favors the opposite gender compared to GPT-4o but is closer to the fairness threshold. Claude-3.5-Sonnet exhibits moderate bias toward males. Notably, these scores are considerably lower than those from template-based probing.
Religious Bias: The bias scores for religion in Figure 4 are comparatively closer among all models. GPT-4o and Llama-3 both exhibit negative bias scores, suggesting some level of bias towards Muslims. GPT-4o exhibits the highest level of bias.
We hypothesize that, the reason for not showing substantial bias in naturally probed examples can be attributed to two points: (1) When a Bangla prompt is provided with a broader and naturally occurring context, the LLMs tend to focus on the overall meaning of the scenario rather than isolating specific characters and attributing gender or religious identities to them. This reduces the likelihood of bias being explicitly reflected in the responses. (2) The guard-rails used in LLMs work better in a natural probing setting.
Key Take-away: The study reveals significant biases in multilingual large language models (LLMs) when generating outputs in Bangla. Gender and religious biases are evident, varying in degree and direction depending on the model and probing method. Template-based probing shows more pronounced biases as opposed to naturally sourced probing.
7 Conclusion
To summarize, our study investigates gender and religious bias in multilingual LLMs within the context of Bangla, utilizing two distinct probing techniques and datasets. The results reveal varying degrees of bias across models and underscore the need for effective debiasing techniques to ensure the ethical use of LLMs in sensitive Bangla-language applications. Additionally, the findings highlight the importance of developing linguistically and culturally aware frameworks for bias measurement. Future research could focus on expanding the dataset to include non-binary genders, additional religious groups, and nuanced sociocultural contexts to better capture the diversity of Bangla-speaking regions.
Limitations
Our study utilized closed-source models like GPT-3.5-Turbo, GPT-4o and Claude-3.5-Sonnet which present reproducibility challenges as they can be updated at any time, potentially altering responses regardless of temperature or top-p settings. We also attempted to conduct experiments with other state-of-the-art models such as Mistral-7b-Instruct 777mistralai/Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Llama-2-7b-chat-hf 888meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023), and OdiaGenAI-BanglaLlama  [^9] OdiaGenAI/odiagenAI-bengali-base-model-v1 (Parida et al., 2023). However, these efforts were hindered by frequent hallucinations and an inability to produce coherent and presentable results. This issue underscores a broader challenge: the current limitations of LLMs in processing Bangla, a low-resource language, indicating a need for more focused development and training on Bangla-specific datasets.
Another limitation of our study is the constrained template based probing, where there is more scope of expansion. Real world downstream tasks such as personalized dialogue generation (Zhang et al., 2018), summarization (Hasan et al., 2021, Bhattacharjee et al., 2023), and paraphrasing (Akil et al., 2022) could also be considered for analyzing bias in LLMs for Bangla.
We also acknowledge that our results may vary with different prompt templates and datasets, constraining the generalizability of our findings. Stereotypes are likely to differ based on the context of the input and instructions. Finally our techniques all utilizes binary identities(male-female, Hindu-Muslim) for the constraints on dataset and techniques used (Please refer to appendix A). Despite these limitations, we believe our study provides essential groundwork for further exploration of social stereotypes in the context of Bangla for LLMs.
Ethical Considerations
Our study focuses on binary gender due to data constraints and existing literature frameworks. We acknowledge the existence of non-binary identities and recommend future research to explore these dimensions for a more inclusive analysis. The same goes for religion. We acknowledge the existence of many other religions in the Bangla-speaking regions, but we focused on the two main religion communities of this ethnolinguistic community.
We acknowledge the inclusion of data points in our dataset that many may find offensive. Since these data are all produced from social media comments, we did not exclude them to reflect real-world social media interactions accurately. This approach ensures our findings are realistic and relevant, highlighting the need for LLMs to effectively handle harmful content. Addressing such language is crucial for developing AI that promotes safer and more respectful online environments.
Appendix
Appendix A Frequency Analysis of Gender and Religion Terms in Two Bangla Corpora
We have kept our studies limited to binary genders and the major religions in Bangla speaking regions. In this section, we provide a quantitative analysis of two major Bangla corpora regarding the frequency distribution of gender and religion realted entities. We show the results in Figure 5.
We extracted the gender and religion related entities from two large corpora, BnWiki [^10] The latest bangla wiki dump used from https://dumps.wikimedia.org/bnwiki/20240901/ and Bangla2B+ Bhattacharjee et al. (2022). It is evident that there is a significant absence of non-binary genders in Bangla. For the male and female words, we used the most common male and female terms in Bangla and later aggregated the results under Men and Women terms in the data showed. The word percentages for transgenders and homosexuals are less than 2%. Note that, we used the term Hijra [^11] https://en.wikipedia.org/wiki/Hijra_(South_Asia) as an umbrella term for non-binary genders, as this semantics is prevalent in South Asia.
Figure 5: Frequency Analysis of Gender and Religious Identities in two large Bangla corpora: BnWiki and Bangla2B+
For the religion related terms, we composed the common religious identity based words in Bangla speaking regions and accommodated for their variations. In both the corpora, we can see that Hindu and Muslim related religious identities comprise of more than 70% of the total identities. Hence considering the availability of dataset, our probing techniques and corpus frequency distribution, we limited our study to binary genders and most common religions.
Appendix B Evaluation Metric Justification
Various metrics have been proposed to evaluate the fairness of LLMs. Disparate Impact compares the proportion of favorable outcomes for a minority group to a majority group, while Statistical Parity compares the percentage of favorable outcomes for monitored groups to reference groups. Metrics such as Equalized Opportunity and Equalized Odds considers ground truth. Since our dataset contains no ground truth, we chose Disparate Impact to evaluate the model responses for binary identities.
In task Q, for category a with outcomes x and y, DI is calculated as:
$DI_{Q}(a)=\frac{P(Q=x|a)}{P(Q=y|a)}$
Since we do not have probability distributions in our case, we use the occurrence frequency of each category instead. However, plotting the graphs with the above formula can be challenging because the values lie in the interval $[0,+\infty)$ with the center line in 1. For an LLM, $DI_{Q}(a)=1$ signifies perfect fairness, while values approaching 00 or +$\infty$+ ∞ indicate extreme bias towards one identity. For example, if $P(Q=female|Gender)=0.01$ and $P(Q=male|Gender)=0.99$, then $DI_{Gender}=\frac{0.01}{0.99}=0.01010101$. Conversely, if $P(Q=female|Gender)=0.99$ and $P(Q=male|Gender)=0.01$, then $DI_{Gender}=\frac{0.99}{0.01}=99$. Though both results reflect significant bias, visually interpreting these results on a graph can be difficult due to the disproportionate scaling.
To address this, we modified the metric as follows:
$Bias\>Score=DI_{Q}(a)=\tanh\left(\log\frac{C_{x}(a)}{C_{y}(a)}\right)$
Here, $C_{z}$ represents the frequency of class z. By applying the logarithmic function, we scale the values proportionally for better interpretation, and we utilize the $\tanh$ function to normalize the bias scores within the interval $[-1,1]$. A $Bias\>Score$ close to 0 indicates fairness, whereas values closer to $-1$ or $1$ indicates extreme bias towards one group or the other.
Appendix C Data Filtration for Naturally Sourced Sentences
The selection criteria for the Explicit Bias Evaluation(EBE) dataset are based on ensuring meaningful and contextually accurate sentences that are neutral from the perspective of gender and religion. In the original BIBED dataset Das et al. (2023), authors created pair for each sentence by replacing the identifying subject, either male-female (for gender) or Hindu-Muslim (for religion) with their respective counterparts (shown in Figure 7). However, in the EBE data, there are many generated pair sentences that are semantically inconsistent for the pair subject as illustrated in the first two columns of Figure 8.
Therefore, for our purpose we refined the dataset and only selected those sentences that are equally probable for either both Male/Female genders and both Hindu/Muslim religion. In order to do that, we prompted GPT-3.5-Turbo to check if the pair sentence of the root sentence is semantically consistent. If altering the gender or religion rendered the sentences factually incorrect or nonsensical, we rejected those as depicted in Figure 8. For instance, sentences involving specific historical figures or roles explicitly or implicitly linked to a particular gender or religion were excluded. The goal was to maintain the integrity of context-specific information, such as unique cultural, historical, or biological aspects, which would be distorted by changing the gender or religion. This approach ensures that the dataset reflects accurate evaluations and free from gender or religion specific information before prompting the models.
Appendix D Annotator’s Agreement on Naturally Selected Data
The final dataset used for naturally sourced probing contains 2416 data points for gender and 1535 data points for religion. Both authors of this paper, being native Bangla speakers, served as annotators. To assess the inter-rater reliability, we utilizied Cohen’s Kappa coefficient, $\kappa$ on a smaller sample (200 for gender and 125 for religion) of the original dataset. We define the following terms: $True\ Positives$ (TP) as the number of samples both annotators selected, $True\ Negatives$ (TN) as the samples both rejected, $False\ Positives$ (FP) as the samples where the first annotator selected but the second rejected, and $False\ Negatives$ (FN) as the samples where the first annotator rejected but the second selected. Details for both sampled dataset is shown in Table 3.
Sampled Gender Dataset (200 data-points)
A1 Selected A1 Rejected
A2 Selected 183 (TP) 3 (FP)
A2 Rejected 4 (FN) 10 (TN)
Sampled Religion Dataset (125 data-points)
A1 Selected A1 Rejected
A2 Selected 115 (TP) 2 (FP)
A2 Rejected 3 (FN) 5 (TN)
Table 3: Binary Classification Confusion Matrix for Annotators’ Agreement
Cohen’s $\kappa$ is a robust statistic used to measure the agreement between two raters who each classify $N$ items into $C$ mutually exclusive categories. Since our dataset involves binary classification (male-female or Hindu-Muslim), we applied a confusion matrix for binary classification and calculated the value of $\kappa$ as follows:
$\kappa=\frac{p_{0}-p_{e}}{1-p_{e}}$
Here, $p_{0}$ represents the observed agreement between the raters and $p_{e}$ refers to the expected agreement due to chance. The probabilities for selecting and rejecting a data point at random are denoted as $p_{1}$ and $p_{2}$, respectively, leading to the following equations:
$\begin{split}p_{0}&=\frac{TP+TN}{N}\\ p_{1}&=\frac{(TP+FN)*(TP+FP)}{N^{2}}\\ p_{2}&=\frac{(TN+FN)*(TN+FP)}{N^{2}}\\ p_{e}&=p_{1}+p_{2}\end{split}$
Based on our smaller sampled dataset, we obtained $\kappa=0.722$ for gender and $\kappa=0.645$ for religion, both indicating substantial agreement between the annotators, thereby confirming the reliability of our dataset.
Appendix E Dataset Statistics
For template based probing, we utilized different categorical adjective words for both gender and religion role prediction as shown in Table 4.
For naturally sourced probing, the average sentence length for Gender topic is 23 words and for Religion topic is 20 words.
Topic Category
Positive
Trait
Negative
Trait
Gender Personality 16 15
Outlook 8 11
Communal 11 9
Religion Ideology 13 16
Outlook 8 7
Occupation 14
Table 4: Count of adjective words used as placeholders for prompt creation
Appendix F Prompt Template
Table 5: The prompt template and an example of prompt for gender role prediction (Note that the translations are only for understanding and not used in prompting). Please note that the translation is not an exact translation of the question. More appropriate translation could have been "He/she is a modest person". But that would have been misleading due to the inclusion of gendered pronouns in English translation, but in fact pronouns in Bangla are gender neutral.
Figure 6: Categories of Adjective words used for templates
Figure 7: Naturally Sourced (EBE) Sentences Examples for Religion and Gender Bias Prediction
Figure 8: Examples of Rejected Sentence and Reason for Rejection
Figure 9: Prompt templates for Bias in Religion and Gender Role Prediction for template based probing. (Note the translations for Gender category. We used ’He/She’ to define the subject in the translations, which could give a false impression of the actual Bangla text. The pronouns in Bangla are gender neutral. But to maintain correspondence and represent first person singular subject in English, we used He/She in the place of subject for English translation. The Bangla sentences are kept neutral, which was used to prompt the model.)